# A note on quasi-perfect morphisms

### 10 Mar 2026 v2

@@ Abstract.
This note records two results concerning quasi-perfect morphisms between Noetherian algebraic spaces. First, we give a new characterization of regular Noetherian algebraic spaces as those for which blowups at closed points are quasi-perfect. Secondly, we study the local behavior of quasi-perfect proper morphisms. In particular, we show that quasi-perfectness of a proper morphism can be detected at their étale local rings, completions, and (strict) Henselizations. As a corollary, the locus of points where a proper morphism is quasi-perfect is Zariski open.
@@
Derived category, quasi-perfect morphism, blowup, regularity, generation (for triangulated categories)
@@ 1. Introduction
Given a quasi-compact quasi-separated morphism $f\colon Y\to X$ of schemes, Neeman’s Brown Representability Theorem provides a right adjoint $f^{\times}$ to the derived pushforward $\mathbf{R}f_{\ast}$ on $D_{\operatorname{qc}}$. While $f^{\times}$ preserves small coproducts for complexes that are uniformly homologically bounded below, this property fails in general. Motivated by [VER69, Lemma 1 & Corollary 2], Lipman–Neeman [LN07] introduced the notion of quasi-perfect morphisms for quasi-compact quasi-separated schemes to remedy this. The defining property is that $f^{\times}$ preserves small coproducts on all of $D_{\operatorname{qc}}$ or, equivalently, that $\mathbf{R}f_{\ast}$ preserve perfect complexes; see [LN07, Theorem 1.2 & Proposition 2.1]. Whilst there are numerous examples of quasi-perfect morphisms—for instance, any proper morphisms of finite Tor dimension between Noetherian schemes—quasi-perfectness generally fails. For example, this is the case for closed immersions into a non-regular Noetherian scheme.
@@
In this short note we make two novel observations regarding quasi-perfect morphisms between algebraic spaces. The first observation is how one can use these morphisms to characterize regularity.
Proposition (Proposition˜3.3).
To the best of our knowledge, this characterization has not appeared before in the literature even for schemes. The simple, but crucial, observation is that a closed point on a Noetherian algebraic space is regular if and only if the blowup at that point is quasi-perfect. This follows from the fact that perfectness of the skyscraper sheaf at a closed point detects regularity of that point; we prove this in Lemma˜3.1, which seems to not have been written down for algebraic spaces before and might be of independent interest. Furthermore, as an application of Proposition˜3.3, we give a new proof that regularity descends along fpqc covers (see Corollary˜4.5).
The second observation concerns the local behavior of quasi-perfectness for proper morphisms. This is summarized in the following result.
Theorem (Theorem˜4.8).
@@
Theorem˜4.8 provides a mechanism for detecting quasi-perfectness by reducing the question to targets that are local rings. The proof relies on several auxiliary lemmas, potentially of independent interest, which allow quasi-perfectness to be checked using specific classes of flat covers (see Lemmas˜4.2 and 4.3). As a corollary, we show that the locus of points $p$, for which the base change to one of the classes of local rings $A_{p}$ from Theorem˜4.8 is quasi-perfect, forms a Zariski open subset.
Corollary (Corollary˜4.9).
Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose that $A_{\#}$ and $s\colon U\to X$ are as in Theorem˜4.8. Then the locus
@@
One can think of $S$ as the ‘the quasi-perfect locus of $f$’. The fact it is always open for proper morphisms appears to be new even for schemes. As an application, we show that the property that the regular locus contains a nonempty open subset (i.e. that $X$ is JJ-0) descends along proper surjective morphisms under mild hypotheses, which are satisfied if, for example, the target is reduced (see Proposition˜4.11).
Finally, we comment on the restriction to proper morphisms in Theorem˜4.8. In view of Proposition˜3.3, this is mild since the morphisms appearing there are proper. In fact, for a separated morphism of finite type between Noetherian schemes, quasi-perfectness already implies properness (see [NEE21, Proposition 0.20]); so this is quite a mild restriction. The properness assumption is required in our arguments to employ local-to-global techniques involving bounded coherent objects (e.g. [AIL10, Theorem 4.1]).
We expect the more general results to be useful in future work on singularities of algebraic spaces, a topic that has recently been gaining traction. For example, recently the minimal model program has been established for excellent algebraic spaces in characteristic zero [LM22].
@@ Acknowledgments.
De Deyn was supported by ERC Consolidator Grant 101001227 (MMiMMa). Lank was supported under the ERC Advanced Grant 101095900-TriCatApp. The authors thank Leovigildo Alonso Tarrío, Fei Peng, and Karl Schwede for looking at an earlier version of this work. Lank would like to dedicate this work to Boz, whose joyous and warm memories left behind will be forever remembered.
@@ 2.2. Algebraic Spaces
Modules and derived categories
@@
Lastly, for future reference, let us note the following lemma which shows that one can check perfectness flat locally; see e.g. [HR17, Lemma 4.1] or [DLR+25, Lemma 2.2]. For the definition of the the fpqc topology, see [STA26, Tag 03MP]; important examples of fpqc covers are fppf, and hence, also smooth and étale covers.
@@
When $X$ is a Noetherian algebraic space, it follows from [DLR+25, Theorem 3.7] that regularity of $X$ is equivalent to $D^{b}_{\operatorname{coh}}(X)=\operatorname{Perf}(X)$. Furthermore, when $X$ is a separated Noetherian algebraic space of finite Krull dimension, by [DLR+25, Theorem 5.1], regularity of $X$ is equivalent to $\operatorname{Perf}(X)$ being strongly generated, i.e. $\operatorname{Perf}(X)=\langle G\rangle_{n+1}$ for some $G\in\operatorname{Perf}(X)$ and $n\geq 0$.
@@ 2.3. Morphisms
We say a morphism of integral algebraic spaces $f\colon Y\to X$ is
  • a modification if it is proper and birational (i.e. generically an isomorphism), see e.g. [STA26, Tag 0ACU] for some characterizations and
  • an alteration if it is proper and generically finite, see e.g. [STA26, Tag 0AD7] for some characterizations.
Usually, in this work the target (and hence, source) of the morphism will be Noetherian. Note that, for either type, the morphism above is, in fact, surjective. This follows as they are by definition dominant (this is implicit in generically finite/an isomorphism) and a dominant proper morphism is surjective. The modifications we will be interested in below arise as blowups. See [STA26, Tag 085P] for some background on blowups of algebraic spaces.
@@
3. Quasi-perfect blowups detect regularity
This section shows a relationship between regularity and having ‘enough’ quasi-perfect ‘probing morphisms’. The following should be compared to [STA26, Tag 00OC] and gives a ‘spacy’ version of this.
Lemma 3.1.
Proof.
[…]
Next, using flat base change, one sees
Lemma 3.2.
Let $X$ be Noetherian algebraic space $X$. If the blowup of $X$ along a closed point $p\in|X|$ induces a quasi-perfect morphisms, then $p$ is regular.
Proof.
[…]
where $f$ is the blowup of $X$ at $p$ and $E$ is the exceptional divisor. By definition of blowing up the exceptional divisor $E$ is an effective Cartier divisor, and so it follows from the distinguished triangle
that $j_{\ast}\mathcal{O}_{E}$ is perfect. Hence, as $f$ is quasi-perfect by assumption, $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}$ is perfect too. To conclude, observe from the commutativity of Diagram˜3.1 that
$0$ $\not=\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}$
$\cong i_{\ast}\mathbf{R}g_{\ast}\mathcal{O}_{E}$
Hence $i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}$ is a summand of $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}\in\operatorname{Perf}(X)$ and so is also perfect. Thus, $p$ is regular by Lemma˜3.1, and we are done. ∎
Proposition 3.3.
Proof.
If $X$ is regular, then for any proper $f$ we have
Hence, any blowup is quasi-perfect as these are proper. The converse follows from Lemma˜3.2 as it suffices to check regularity at closed points (see e.g. [DLR+25, Lemma 3.2]). ∎
Lemma 3.4.
Proof.
Corollary 3.5.
Proof.
This follows immediately from Lemma˜3.2. Indeed, the blowups appearing in the proof have integral target and so are modifications by Lemma˜3.4. Moreover, as modifications are alterations the case of alterations also follows. ∎
Remark 3.6.
In the special case where there exists an alteration or modification $f\colon Y\to X$ of algebraic spaces where $Y$ is regular and $f$ quasi-perfect, then one can show $X$ regular more easily using ‘proper descent’. For example, this follows from [HLL+25, Corollary 6.3], and the fact that $D^{b}_{\operatorname{coh}}=\operatorname{Perf}$ implies regularity. So, the novelty of Corollary˜3.5 lies in its independence of the existence of such a morphism, giving a different characterization of regularity.
4. On proper quasi-perfect morphisms
Here we study quasi-perfect morphisms which are proper. We start with a few lemmas that allow one to check quasi-perfectness using flat covers.
@@ 4. On proper quasi-perfect morphisms
Hence, Lemma˜4.1 implies $f^{\prime}$ is a quasi-perfect morphism.
For the second claim, assume now that $f$ is surjective and $f^{\prime}$ is quasi-perfect. Then $\mathbf{R}f^{\prime}_{\ast}(s^{\prime})^{\ast}Q$ is perfect and so is $s^{\ast}\mathbf{R}f_{\ast}Q$ by Equation˜4.2. As $s$ is surjective, so is $s^{\prime}$, and so it is an fpqc cover as it is quasi-compact. Therefore, $\mathbf{R}f_{\ast}Q$ is perfect as this can be checked flat locally by Lemma˜2.1. By Lemma˜4.1, we have that $f$ is quasi-perfect, which completes the proof. ∎
The following is a variation on the previous.
Lemma 4.3.
Let $f\colon Y\to X$ be a morphism of quasi-compact quasi-separated algebraic spaces with $X$ an affine scheme. Suppose $s\colon U\to X$ is a flat surjective morphism from an affine scheme. Then $f$ is quasi-perfect if and only if the base change of $f$ along $s$ is quasi-perfect.
Proof.
The forward direction follows from Lemma˜4.2. For the converse, denote the base change of $f$ along $s$ as in Diagram˜4.1. Again, choose a compact generator $Q$ of $Y$. By assumption $f^{\prime}$ is quasi-perfect, hence, using Equation˜4.2, $s^{\ast}\mathbf{R}f_{\ast}Q$ is perfect. However, as $s$ corresponds to a faithfully flat ring map, it follows from [STA26, Tag 068T] that $\mathbf{R}f_{\ast}Q$ is perfect. Thus, Lemma˜4.1 shows that $f$ is quasi-perfect. ∎
Using Lemma˜4.2, we obtain an alternative proof of the fact that regularity can be checked fpqc locally. Before doing so, we recall a fact whose proof we add for lack of finding a reference.
Lemma 4.4.
Proof.
Corollary 4.5.
Proof.
First, as $X$ is quasi-compact, we can reduce to an fpqc cover with finitely many components (i.e. $|I|<\infty$). Secondly, if needed, we can refine the cover so that each $U_{i}$ is an affine scheme (e.g. take étale surjective morphisms from affine schemes). Next, consider the induced morphism $f\colon\bigsqcup_{i\in I}U_{i}\to X$ which is faithfully flat by construction and moreover is quasi-affine by Lemma˜4.4 as $U:=\bigsqcup_{i\in I}U_{i}$ is (isomorphic to) an affine scheme. To show $X$ is regular, it suffices by Lemma˜3.2 to show that blowups along nonzero ideal sheaves are quasi-perfect. However, as blowups commute with flat base change, see e.g. [STA26, Tag 0805], we have that the base change to $U$ is again such a blow-up. Consequently, as $U$ is regular, it is quasi-perfect. Thus, the original blow-up of $X$ is also quasi-perfect, by Lemma˜4.2, showing the claim. ∎
Remark 4.6.
There are three important cases of faithfully flat morphisms that we will apply Lemma˜4.3.
@@
We now move to our next main result.
@@
The proof of ($\star$4) is similar to that of Lemma˜4.2; however, note that $\operatorname{Spec}(\mathcal{O}_{U,p})\to X$ is generally not surjective so we need to do some extra work. Pick a point $p\in U$ and look at the fibered square
@@
Again, choose any compact generator $Q$ of $Y$. The hypothesis implies that $\mathbf{R}(f_{p})_{\ast}t^{\ast}_{p}Q$ is perfect, and so flat base change tells us $s_{p}^{\ast}\mathbf{R}f_{\ast}Q=(s^{\ast}\mathbf{R}f_{\ast}Q)_{p}$ is perfect. As $U$ is Noetherian, and as $p\in U$ was arbitrary, we have that $s^{\ast}\mathbf{R}f_{\ast}Q$ is perfect (this can be checked affine locally via [AIL10, Theorem 4.1])–this uses the fact that $s^{\ast}\mathbf{R}f_{\ast}Q$ is bounded and coherent by virtue of $f$ being proper. Consequently, so is $\mathbf{R}f_{\ast}Q$ by Lemma˜2.1 and $f$ is quasi-perfect by Lemma˜4.1. ∎
@@
Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose that $A_{\#}$ and $s\colon U\to X$ are as in Theorem˜4.8. Then the locus
@@
By the proof of Theorem˜4.8 (i.e. using Lemma˜4.3) we know that
@@
To this end, observe by Equation˜2.1 that $\mathcal{O}_{X,s(u)}^{sh}=\mathcal{O}^{sh}_{U,u}$ for any $u\in U$. As $\mathcal{O}_{U,u}\to\mathcal{O}^{sh}_{U,u}$ is faithfully flat, quasi-perfectness of base change to either of them is equivalent by Lemma˜4.3. Therefore,
@@
As a consequence of Equation˜4.3 and that $s\colon U\to|X|$ is an open quotient map, it follows that $S_{X}$ is open if and only if $S_{U}$ is open. Thus, picking another cover if necessary we may assume $U$ is affine and it suffices to show $S_{U}$ is open to finish the proof.
@@
One can think of the subset $S$ appearing in Corollary˜4.9 as the ‘quasi-perfect locus of $f$’. It follows from the first part of the proof that in general, i.e. regardless of properness of $f$, one can compute it étale locally (Equation˜4.3) and that the specific choice of $A_{\#}$ does not matter. Moreover, this subset is open when $f$ is proper; it would be interesting to know whether or not this holds for more general morphisms.
@@
Let $X$ be a Noetherian algebraic space containing at least one codimension 0 point [^1] See [STA26, Tag 04NA] for the relevant definition. whose étale local ring is reduced (e.g. this is the case when $X$ is reduced). Suppose further that $f\colon Y\to X$ is a proper surjective morphism. Then the regular locus of $X$ contains a nonempty open subset (i.e. is JJ-0) whenever $Y$ has this property.
@@
  • Next, let $S$ denote the open subset from Corollary˜4.9; as $\eta\in S$, since $\mathcal{O}_{X,\eta}$ is a field by assumption, $S$ is non-empty. It follows from Theorem˜4.8 that the restriction $f^{-1}(S)\to S$ is quasi-perfect. So we may assume $f$ is quasi-perfect.
@@
In order to finish the proof, is suffices to show that $X$ is regular. This follows from ‘proper descent’, as already indicated in Remark˜3.6; but we give a self-contained proof, more in line with that of Lemma˜3.2, for the convenience of the reader.
So, let $p\in X$ be a closed point and denote the associated closed immersion by $i\colon\mathrm{Spec}(\kappa(p))\to X$. Consider the fibered square

### 17 Mar 2026 v3

@@ Abstract.
This note is concerned with quasi-perfect morphisms between Noetherian algebraic spaces. In particular, we study the local behavior of quasi-perfect proper morphisms. We show that quasi-perfectness of a proper morphism can be detected at the étale local rings of points of the target, as well as their completions and (strict) Henselizations. As a corollary, we obtain that the locus of points where a proper morphism is quasi-perfect is Zariski open.
@@
Algebraic spaces, quasi-perfect morphism, generation (for triangulated categories), regularity
@@ 1. Introduction
Given a quasi-compact and quasi-separated morphism $f\colon Y\to X$ of schemes, Neeman’s Brown Representability Theorem provides a right adjoint $f^{\times}$ to the derived pushforward $\mathbf{R}f_{\ast}$ on $D_{\operatorname{qc}}$, the derived category of complexes with quasi-coherent cohomology. While $f^{\times}$ preserves small coproducts for complexes that are uniformly homologically bounded below, this property fails in general. Motivated by [VER69, Lemma 1 & Corollary 2], Lipman–Neeman [LN07] introduced the notion of a quasi-perfect morphism to remedy this.
The defining property is that $f^{\times}$ preserves small coproducts on all of $D_{\operatorname{qc}}$ which is equivalent to $\mathbf{R}f_{\ast}$ preserving perfect complexes; see [LN07, Theorem 1.2 & Proposition 2.1]. There are numerous examples of quasi-perfect morphisms, e.g. any proper morphisms of finite Tor dimension between Noetherian schemes. However, quasi-perfectness generally fails; for example, for closed immersions into nonregular Noetherian schemes.
@@
In this note, we study the local behavior of quasi-perfectness for proper morphisms between algebraic spaces.
Theorem (Theorem˜3.4).
@@
To the best of our knowledge, the observed behavior is new even for schemes. Theorem˜3.4 provides a mechanism for detecting quasi-perfectness by reducing the question to targets that are local rings. The proof relies on several auxiliary lemmas, potentially of independent interest, that allow for quasi-perfectness to be checked using specific classes of flat covers (see Lemmas˜3.2 and 3.2).
As a corollary, we have the following.
Corollary (Corollary˜3.6).
Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose that $A_{\#}$ and $s\colon U\to X$ are as in Theorem˜3.4. Then the locus
@@
One can think of the open subset $S$ as the ‘the quasi-perfect locus of $f$. The fact it is always open for proper morphisms appears to be new even for schemes. As an application, we show the following.
Proposition.
Let $X$ be a reduced Noetherian algebraic space. Suppose further that $f\colon Y\to X$ is a proper surjective morphism. Then the regular locus of $X$ contains a nonempty open subset (i.e. is $J\textrm{-}0$) whenever $Y$ has this property.
Again, to the best of our knowledge, this is new even in the case of schemes. In fact, we prove a slightly more general statement, see Proposition˜3.8. Furthermore, in Appendix˜A, we record a characterization for regularity of Noetherian algebraic spaces in terms of quasi-perfectness of blowups.
Finally, we comment on the restriction to proper morphisms in Theorem˜3.4. In practice, it is a mild restriction. Indeed, for a separated morphism of finite type between Noetherian schemes, quasi-perfectness already implies properness (see [NEE21, Proposition 0.20]). The properness assumption is required in our arguments to employ local-to-global techniques involving bounded coherent objects (e.g. [AIL10, Theorem 4.1]).
We expect the more general results to be useful in future work on singularities of algebraic spaces. This a topic that has recently been gaining traction. For example, recently the minimal model program has been established for excellent algebraic $\mathbb{Q}$-spaces [LM22].
@@ Acknowledgments.
De Deyn was supported by ERC Consolidator Grant 101001227 (MMiMMa). Lank was supported under the ERC Advanced Grant 101095900-TriCatApp. The authors thank Leovigildo Alonso Tarrío, Bhargav Bhatt, Linquan Ma, Fei Peng, and Karl Schwede for discussions and/or comments. Lank would like to dedicate this work to Boz, whose joyous and warm memories left behind will be forever remembered.
@@ 2.2. Algebraic Spaces
Derived categories and functors
@@
Lastly, for future reference, let us note the following lemma which shows that one can check perfectness flat locally; see e.g. [HR17, Lemma 4.1] or [DLM+25, Lemma 2.2]. For the definition of the the fpqc topology, see [STA26, Tag 03MP]; important examples of fpqc covers are fppf, and hence, also smooth and étale covers.
@@
When $X$ is a Noetherian algebraic space, it follows from [DLM+25, Theorem 3.7] that regularity of $X$ is equivalent to $D^{b}_{\operatorname{coh}}(X)=\operatorname{Perf}(X)$. Furthermore, when $X$ is a separated Noetherian algebraic space of finite Krull dimension, by [DLM+25, Theorem 5.1], regularity of $X$ is equivalent to $\operatorname{Perf}(X)$ being strongly generated, i.e. $\operatorname{Perf}(X)=\langle G\rangle_{n+1}$ for some $G\in\operatorname{Perf}(X)$ and $n\geq 0$.
@@ 2.3. Morphisms
3. Proofs
We prove our main results. To start, we need a few lemmas that allow one to check quasi-perfectness using flat covers.
@@ 3. Proofs
Hence, Lemma˜3.1 implies $f^{\prime}$ is a quasi-perfect morphism.
For the second claim, assume now that $s$ is surjective and $f^{\prime}$ is quasi-perfect. Then $\mathbf{R}f^{\prime}_{\ast}(s^{\prime})^{\ast}Q$ is perfect and so is $s^{\ast}\mathbf{R}f_{\ast}Q$ by Equation˜3.2. As $s$ is surjective, so is $s^{\prime}$, and so it is an fpqc cover as it is quasi-compact. Therefore, $\mathbf{R}f_{\ast}Q$ is perfect as this can be checked flat locally by Lemma˜2.1. By Lemma˜3.1, we have that $f$ is quasi-perfect, which completes the proof. ∎
There are three important cases of faithfully flat morphisms that we will apply Lemma˜3.2.
@@
We now move to our main result.
@@
Before proving this, we recall a fact whose proof we add for lack of finding a reference.
Lemma 3.5.
@@
The proof of ($\star$3) is similar to that of Lemma˜3.2; however, note that $\operatorname{Spec}(\mathcal{O}_{U,p})\to X$ is generally not surjective so we need to do some extra work. Pick a point $p\in U$ and look at the fibered square
@@
Again, choose any compact generator $Q$ of $Y$. The hypothesis implies that $\mathbf{R}(f_{p})_{\ast}t^{\ast}_{p}Q$ is perfect, and so flat base change tells us $s_{p}^{\ast}\mathbf{R}f_{\ast}Q=(s^{\ast}\mathbf{R}f_{\ast}Q)_{p}$ is perfect. As $U$ is Noetherian, and as $p\in U$ was arbitrary, we have that $s^{\ast}\mathbf{R}f_{\ast}Q$ is perfect (this can be checked affine locally via [AIL10, Theorem 4.1])–this uses the fact that $s^{\ast}\mathbf{R}f_{\ast}Q$ is bounded and coherent by virtue of $f$ being proper. Consequently, so is $\mathbf{R}f_{\ast}Q$ by Lemma˜2.1 and $f$ is quasi-perfect by Lemma˜3.1. ∎
@@
Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose that $A_{\#}$ and $s\colon U\to X$ are as in Theorem˜3.4. Then the locus
@@
By the proof of Theorem˜3.4 (i.e. using Lemma˜3.2) we know that
@@
To this end, observe by Equation˜2.1 that $\mathcal{O}_{X,s(u)}^{sh}=\mathcal{O}^{sh}_{U,u}$ for any $u\in U$. As $\mathcal{O}_{U,u}\to\mathcal{O}^{sh}_{U,u}$ is faithfully flat, quasi-perfectness of base change to either of them is equivalent by Lemma˜3.2. Therefore,
@@
As a consequence of Equation˜3.3 and that $s\colon U\to|X|$ is an open quotient map, it follows that $S_{X}$ is open if and only if $S_{U}$ is open. Thus, picking another cover if necessary we may assume $U$ is affine and it suffices to show $S_{U}$ is open to finish the proof.
@@
One can think of the subset $S$ appearing in Corollary˜3.6 as the ‘quasi-perfect locus of $f$’. It follows from the first part of the proof that in general, i.e. regardless of properness of $f$, one can compute it étale locally (Equation˜3.3) and that the specific choice of $A_{\#}$ does not matter. Moreover, this subset is open when $f$ is proper; it would be interesting to know whether or not this holds for more general morphisms.
@@
Let $X$ be a Noetherian algebraic space containing at least one codimension 0 point [^1] See [STA26, Tag 04NA] for the relevant definition. whose étale local ring is reduced (e.g. this is the case when $X$ is reduced). Suppose further that $f\colon Y\to X$ is a proper surjective morphism. Then the regular locus of $X$ contains a nonempty open subset (i.e. is $J\textrm{-}0$) whenever $Y$ has this property.
@@
  • Next, let $S$ denote the open subset from Corollary˜3.6; as $\eta\in S$, since $\mathcal{O}_{X,\eta}$ is a field by assumption, $S$ is non-empty. It follows from Theorem˜3.4 that the restriction $f^{-1}(S)\to S$ is quasi-perfect. So we may assume $f$ is quasi-perfect.
@@
In order to finish the proof, is suffices to show that $X$ is regular. So, let $p\in X$ be a closed point and denote the associated closed immersion by $i\colon\mathrm{Spec}(\kappa(p))\to X$. Consider the fibered square
@@
Appendix A Quasi-perfect blowups detect regularity
The goal of this appendix is to prove Proposition˜A.3, which gives a relationship between regularity of a Noetherian algebraic space $X$ and having ‘enough’ quasi-perfectness of morphisms with target $X$. In an earlier version of our work, we recorded Proposition˜A.3 as an observation. However, Bhatt privately communicated to us preliminary work of his, which is to appear later, that subsumes Proposition˜A.3. However, we felt it worthwhile to record this (and some of the lemmas used in its proof) in an appendix.
We start with the following lemma, which should be compared to [STA26, Tag 00OC] and gives a ‘spacy’ version of this.
Lemma A.1.
Proof.
[…]
Next, using flat base change, one sees
The following lemma is the main crucial, albeit simple, observation.
Lemma A.2.
Let $X$ be Noetherian algebraic space $X$. If the blowup of $X$ along a closed point $p\in|X|$ induces a quasi-perfect morphism, then $p$ is regular.
Proof.
[…]
where $f$ is the blowup of $X$ at $p$ and $E$ is the exceptional divisor. By definition of blowing up, the exceptional divisor $E$ is an effective Cartier divisor, and so it follows from the distinguished triangle
that $j_{\ast}\mathcal{O}_{E}$ is perfect. Hence, as $f$ is quasi-perfect by assumption, $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}$ is perfect too. To conclude, observe from the commutativity of Diagram˜A.1 that
$0$ $\not=\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}$
$\cong i_{\ast}\mathbf{R}g_{\ast}\mathcal{O}_{E}$
Hence $i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}$ is a summand of $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}\in\operatorname{Perf}(X)$ and so is also perfect. Thus, $p$ is regular by Lemma˜A.1, and we are done. ∎
Proposition A.3.
Proof.
If $X$ is regular, then for any proper $f$ we have
Hence, any blowup is quasi-perfect as these are proper. The converse follows from Lemma˜A.2 as it suffices to check regularity at closed points (see e.g. [DLM+25, Lemma 3.2]). ∎
Recall that a morphism of integral algebraic spaces $f\colon Y\to X$ is a modification (resp. alteration) if it is proper and birational, i.e. generically an isomorphism (resp. generically finite). See e.g. [STA26, Tag 0ACU] and [STA26, Tag 0AD7] for some characterizations.
Corollary A.4.
Proof.
This follows immediately from Lemma˜A.2. Indeed, the blowups appearing in the proof have integral target and so are modifications by Lemma˜A.5 below. Moreover, as modifications are alterations the case of alterations also follows. ∎
Lemma A.5.
Proof.
To finish, using Lemmas˜3.2 and A.2, we give an alternative proof of the fact that regularity can be checked fpqc locally.
Corollary A.6.
Proof.
First, as $X$ is quasi-compact, we can reduce to an fpqc cover with finitely many components (i.e. $|I|<\infty$). Secondly, if needed, we can refine the cover so that each $U_{i}$ is an affine scheme (e.g. take étale surjective morphisms from affine schemes). Next, consider the induced morphism $f\colon\bigsqcup_{i\in I}U_{i}\to X$ which is faithfully flat by construction and moreover is quasi-affine by Lemma˜3.5 as $U:=\bigsqcup_{i\in I}U_{i}$ is (isomorphic to) an affine scheme. To show $X$ is regular, it suffices by Lemma˜A.2 to show that blowups along nonzero ideal sheaves are quasi-perfect. However, as blowups commute with flat base change, see e.g. [STA26, Tag 0805], we have that the base change to $U$ is again such a blow-up. Consequently, as $U$ is regular, it is quasi-perfect. Thus, the original blow-up of $X$ is also quasi-perfect, by Lemma˜3.2, showing the claim. ∎
Remark A.7.